OpenStudy (pphalke):

Show that each pair is a solution of the given equation. Use the graph of the two points to determine another solution. y=1/2x+1; (6,4) and (-10,-4) is (6,4) a solution of the above vacation is (-10,-4) a solution of the above vacation Use the points given above and the graphing tool to graph the equation. is (6,4) a solution of y=1/2x+1? yes it is -4=1/2(-10)+1 -4=-4 yes is -10,-4 a solution? -4=1/2(-10)+1 -4=-4 yes

OpenStudy (pphalke):

I got everything good need help

OpenStudy (pphalke):

@Aveline

OpenStudy (aveline):

It looks like those are both solutions.

OpenStudy (pphalke):

I need help in making a graph, it gives me one but don't know how the graph should look

OpenStudy (pphalke):

@sbuck98 do u know how to make a graph for this? I'm stuck

OpenStudy (sbuck98):

I'm not any good at math but, @MathMusician @ganeshie8

OpenStudy (aveline):

https://www.desmos.com/calculator Use this

OpenStudy (mathmusician):

(6,4) is not a solution

OpenStudy (mathmusician):

Just plug in the points into the equation (x,y)-->x = 6, y = 4

OpenStudy (mathmusician):

x = -10 y = -4

OpenStudy (aveline):

y=1/2x+1; (6,4) This works 4=6*(1/2)+1

OpenStudy (pphalke):

how do I graph it on the plane

OpenStudy (aveline):